OpenStudy (anonymous):

What is the prime factorization of 225? A. 2 x 3 x 5 x 5 B. 9 x 25 C. 3 x 3 x 5 x 5

jhonyy9 (jhonyy9):

15*15=225

OpenStudy (anonymous):

C

jhonyy9 (jhonyy9):

C yes jimmyrep right

OpenStudy (turingtest):

both B and C multiply to 225: 9x25=225 3x3x5x5=225 but only C has all prime numbers, so that is the PRIME factorization. 9=3x3 so not prime 25=5x5 so not prime

OpenStudy (turingtest):

A is just plain wrong because it doesn't even make 225 2 x 3 x 5 x 5=150

OpenStudy (anonymous):

i dont know if your aware of this but a quick way to find if a number is divisible by 3 is to keep adding its digits auntil you have one digit and it this is 3 6 or 9 then you have a multiple of 3. so 2 + 2 + 5 = 9 - its divisible by 3

OpenStudy (anonymous):

Thank you all! (:

OpenStudy (anonymous):

yw